OpenStudy (anonymous):

If f(x) = x 2 + 1 and g(x) = 3x + 1, find g(20) + f(6).

OpenStudy (whpalmer4):

You should know how to do these — we've told you for a couple of problems in quick succession.

OpenStudy (anonymous):

g(20)=3(20)+1 g(20)=61 f(6)=6^2+1 f(6)=37 g(20)+f(6)=98